# Reputation in the Shadow of Exit

### v4

@@ 4.2 Guaranteeing Stackelberg Payoffs
What classes of reputation results might one hope for in the game with exit? Recall that classical reputation results hold only after many121212“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff.
@@
The first problem, as evidenced by Proposition 2, can be resolved by adding an appropriate incentive compatibility constraint in $(\delta,L(\cdot,\cdot))$. The second problem, however, is not so easy to fix: indeed, I will need the full force of the Markovian refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
@@
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\bar{\delta}<1$ such that for all $\delta>\bar{\delta}$ and each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$\delta|V_{*}(\delta)L(a,c^{*}(\delta))-\bar{V}_{0}(\delta)|>(\geq)\left|v(a,0)-v(a,c^{*}(\delta))L(a,c^{*}(\delta))\right|$
This incentive compatibility condition is stronger than the one found in the folk theorem, Proposition 2. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 2 (see Corollary 1). For now, it delivers the following theorem, (modulo one last definition for the converse).
@@
Theorem 1.
$\limsup_{\delta\to 1}|\underline{U}^{M}(\delta)-V_{*}(\delta)|=0$
for any $\mu_{0}(\omega_{R})<1$. Conversely, if $(f,v,u)$ is monotone and the long-run player does not asymptotically satisfy weak $c^{*}$ incentive compatibility, then for all $\delta<1$,
@@
The forward direction. Suppose the long-run player asymptotically satisfies $c^{*}$ incentive compatibility, and fix $\delta>\bar{\delta}$ in the definition. We show in all Markov perfect equilibria the long-run player plays their commitment action $c^{*}\in c^{*}(\delta)\cap\Omega$ and short-run players best respond, securing the long-run player $V_{*}(\delta)$. To do so, we establish a number of lemmas whose (technical) proofs are relegated to the Proof of Lemma 3
@@
Lemma 3 (along with Proposition 3) guarantees that, off the equilibrium path, the highest payoff the long-run player can obtain is $\bar{V}_{0}(\delta)$, regardless of their discount rate. Our next lemma disciplines the set of actions the long-run player can take in any equilibrium.
@@
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\in\{0,c^{*}(\delta)\}$.
Lemma 4, along with Lemma 3 are enough to imply (following an argument similar to the first part of the proof of Proposition 2 and taking $\alpha=0$), that always playing $c^{*}$ is a Markov perfect equilibrium when $\delta>\bar{\delta}$.
@@
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]<f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0. Finally, we show that the long-run player prefers not to mix in Markovian equilibria.
@@
The Markovian refinement here disciplines short-run player learning; long-run players, in a Markovian equilibrium, begin teaching short-run players about play at time $t=0$ (instead of needing to slowly form their reputation). Reputation thus only has a “backwards looking” component: instead of needing to teach short-run players of future play by taking some commitment action today, they need only maintain their current reputation that they will continue to do what they have already done. In reputational games with exit, this backwards-looking force is the economically important one (i.e. the one that is not blocked by the exit structure).
Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. It is satisfied for example whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply mixed Stackelberg strategies cannot do strictly better than pure ones, which rules out the long-run player from benefiting from the mixed Nash equilibria that can arise when Markovian equilibrium need not be pure (i.e. asymptotic $c^{*}$ incentive compatibility is violated, as per Lemma 5).
@@
Suppose $v(a,c)=v(c)$, $L(a,c)$ is constant in $a$, $|\mathcal{N}(0)|=1$, and $V^{*}(\delta)=V_{*}(\delta)$. Then the following are equivalent.
@@
    There exists $\bar{\mu}_{0}>0$ and $\delta<\bar{\delta}$ such that for all $\mu_{0}<\bar{\mu}_{0}$, $\delta>\bar{\delta}$, $\underline{U}^{M}(\delta)=V^{*}(\delta)$.
@@
Theorem 1 and Corollary 1 are the main theoretical takeaways of this paper. First, they identify a sufficient condition—asymptotic $c^{*}$-incentive compatibility—that negates the effect of exit and ensures the long-run player their Stackelberg payoff in Markovian equilibrium. Moreover, by Proposition 3, the long-run player attains a strictly higher payoff when they have incentives for reputation formation than their best Markovian payoff in the game without reputation. Finally, under a monotonicity condition, these conditions are tight—the long-run player attains their Stackelberg payoff exactly when they satisfy $c^{*}$ incentive compatibility. Corollary 1 then implies that, under some regularity conditions, Markovian equilibria uniquely select the long-run player’s payoff only when such a payoff is feasible, highlighting exit as the key friction hindering reputational incentives in this setting.
4.3 Non-Markovian equilibrium
@@
Theorem 1 focuses on Markovian strategies and leaves open the question of when reputation can help refine payoffs away from the Markovian case. Without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markovian strategies? I next show the answer is no—even with asymptotic $c^{*}$ incentive compatibility, one cannot refine the payoff set in general, as the long-run player’s payoff may be arbitrarily low.
@@
Fix payoffs $(u,v)$ such that there exists $c$ where $\min_{a\in\mathcal{N}(c)}v(a,c)>0$. Then for any $\varepsilon>0$, there exists survival probability function $f$ such that
  1. (1)
    The folk theorem holds.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1,\delta\to 1}\underline{U}_{0}(\delta)\leq\varepsilon$.
The first assumption is a nontriviality condition guaranteeing the existence of a non-Nash commitment posture, conditional on a high probability of survival. The first limit in the second statement guarantees that the probability of commitment is not “so high” that what the long-run player does will not affect their payoffs (i.e. if $\mu_{0}(\omega_{c^{*}})\approx 1$). The second limit guarantees that the folk theorem holds, and shows that when it does reputation effects cannot, in general, refine the set of payoffs.
Proposition 4 highlights the two important roles of Markovian equilibrium in establishing the positive reputation result of Theorem 1. First, it disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward tomorrow actions which are not the commitment action. Second, it ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period (c.f. the discussion succeeding Theorem 1. Markovian beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today. The proof is relegated to the PROOF OF PROPOSITION 4 The table below summarizes the main takeaway from the main results in this section and is a core takeaway of the paper.
non-Markov Markov
Reputation Payoff Set Ambiguous (Proposition 4) Stackelberg Payoff (Theorem 1)
No Reputation Folk Theorem (Proposition 2) Static Nash Payoff (Proposition 3)
@@ 5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how cruelly they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\bar{c}]$ and moreover suppose $\kappa(\bar{c})>1$, so that the government never chooses the maximally feasible value of punishment [^13] Note $C=[0,\bar{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(1,0)$, after which it is sufficient to assume $C=\{0,c^{*}(1,0)\}$ for the remainder of the application. [^14] This is an Inada-type condition that ensures the government’s problem is always well-defined.
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(P_{\theta})=\tau^{2}$ [^15] In practice, I will consider the case where $\tau^{2}$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength.. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfectly signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
@@
    The first-order conditional distribution $F_{\theta}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
@@
    The second-order conditional distribution $F_{x}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals. The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the state is not in a dominance region.
Assumption 3 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable in the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature [^16] I am not aware of a global games setting that does not satisfy these assumptions; for example it nests normal-normal information structures..
Having seen their private signals, the dissidents independently and simultaneously choose in each period an action $a_{i}\in\{0,1\}$, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents defect (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt is successful, and is given by $f(A,c)=\mathbb{P}_{\theta}(A\geq\theta)$ and is independent of $c$.
Player’s payoffs are given by
@@
Proposition 5 implies not only that $\mathcal{N}(c)$ is single-valued, but that both the set of states which successfully guard against attacks and the cutoff state are themselves monotone around a cutoff, whose evolution is well-behaved. These properties will be useful in the proof of 2 and is proven in the PROOF OF PROPOSITION 5
Property (3) along with the existence of dominance regions implies that Assumption 1 is satisfied. Moreover, since $f(a,c)v(a,c)=\mathbb{P}_{\theta}(A\geq\theta)\cdot(1-\kappa(c))$ and $A^{*}(0)=1$ (i.e. all players attack since it is the weakly dominated strategy), Assumption 2 is satisfied as well.
Finally, the game is clearly monotone when $A$ is ordered in the standard way in $\mathbb{R}$, and satisfies the conditions of Corollary 1. Hence the globalized game satisfies the assumptions of the general model.

### v5

@@ 3.2 Guaranteeing Stackelberg Payoffs
What classes of reputation results might one hope for in the game with exit? Recall that classical reputation results hold only after many111111“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff.
@@
The first problem, as evidenced by Proposition 1, can be resolved by adding an appropriate incentive compatibility constraint for the long-run player. The second problem, however, is not so easy to fix: indeed, I will need the full force of the Markov refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
@@
Say the long-run player satisfies (weak) $c^{*}$ incentive compatibility if for each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$(V_{*})L(a,c^{*})-\overline{V}_{0}>(\geq)\left(v(a,0)-v(a,c^{*})L(a,c^{*})\right)$
This incentive compatibility condition is stronger than the one found in the folk theorem, Proposition 1. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise on the equilibrium path, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 1 (see Corollary 1). For now, it delivers the following theorem, (modulo one last definition for the converse).
@@
Theorem 1 (Markov Commitment).
Let $\Omega=\{0,c^{*}\}$ with $\mu_{0}(\omega_{R})<1$. If the long-run player satisfies $c^{*}$ incentive compatibility, then in all Markov equilibrium $\sigma_{L}(h^{t})\equiv c^{*}$. In particular, $\underline{U}^{M}=V_{*}$.
Conversely, if $(f,v,u)$ is monotone and the long-run player does not satisfy weak $c^{*}$ incentive compatibility, then
$\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|>0.$
@@
The forward direction. Suppose the long-run player satisfies $c^{*}$ incentive compatibility. We show in all Markov perfect equilibria the long-run player plays their commitment action $c^{*}$ and short-run players best respond, securing the long-run player $V_{*}$. To do so, we establish a number of lemmas whose (technical) proofs are relegated to the PROOF OF LEMMA 1
@@
Lemma 1 (along with Proposition 2) guarantees that, off the equilibrium path, the highest payoff the long-run player can obtain is $\overline{V}_{0}$. Our next lemma disciplines the set of actions the long-run player can take in any equilibrium.
@@
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\in\{0,c^{*}\}$.
Lemma 2, along with Lemma 1 are enough to imply (following an argument similar to the first part of the proof of Proposition 1 and taking $\alpha=1$), that always playing $c^{*}$ is a Markov perfect equilibrium.
@@
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a,c^{*})[v(a,c^{*})+V_{*}]$
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0. Finally, we show that the long-run player prefers not to mix in Markov equilibria.
@@
The Markov refinement here disciplines short-run player learning; long-run players, in a Markov equilibrium, begin teaching short-run players about play at time $t=0$ (as opposed to needing to slowly form their reputation). Reputation thus only has a “backwards looking” component: instead of needing to teach short-run players about future play by taking some commitment action today, they need only maintain their current reputation that they will continue to do what they have already done. Theorem 1 therefore can be interpreted as saying in reputational games with exit that this backwards-looking force is the economically relevant one, i.e. the one that is not blocked by the exit structure.
Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. For example, it is satisfied whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply that when the long-run player is indifferent between several actions, they cannot secure a higher payoff than their pure Stackelberg payoff. This rules out situations where the long-run player benefits from mixed Nash equilibrium. In particular, it ensures that when mixing occurs in the first period (which is used to prove the converse to Theorem 1), the long-run player must strictly underperform their commitment payoff in such an equilibrium.
@@
Suppose $v(a,c)=v(c)$, $L(a,c)=\mathcal{L}(c)$ for some $\mathcal{L}(c)$, $|\mathcal{N}(0)|=1$, and $V^{*}=V_{*}$. Then the following are equivalent.
@@
    There exists $\overline{\mu}_{0}>0$ such that for all $\mu_{0}<\overline{\mu}_{0}$, $\delta>\overline{\delta}$, $\underline{U}^{M}=V^{*}$.
@@
[…]
3.3 Non-Markov equilibrium
@@
Theorem 1 focuses on Markov strategies and leaves open the question of when reputation can help refine payoffs away from the Markov case. Without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markov strategies? I next show the answer is no—even with $c^{*}$ incentive compatibility, one cannot refine the payoff set in general, as the long-run player’s payoff may be arbitrarily low.
Proposition 3 (Anti-Commitment.).
Fix payoffs $(u,v)$ such that there exists $c$ where
$\min_{a\in\mathcal{N}(c)}v(a,c)>0$. Then for any $\varepsilon>0,M>0$, there exists $f$ such that
  1. (1)
    The game admits a folk theorem and $V^{*}\geq M$.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq\varepsilon$.
The first assumption is a nontriviality condition guaranteeing the existence of a non-Nash commitment posture, conditional on a high probability of survival. The limit in the second statement guarantees the probability of commitment is not “so high” that what the rational long-run player does will not affect their payoffs (i.e. if $\mu_{0}(\omega_{c^{*}})\approx 1$).
Proposition 3 is stronger than the simple statement that the long-run player’s payoff can be arbitrarily low, which can be achieved by simply setting $f(a,c)\leq\eta$ for all pairs $(a,c)$ and arbitrarily low $\eta$. In particular, the first statement guarantees that the long-run player attains a low payoff even if (1) the folk theorem holds, so they can attain their Stackelberg payoff, and (2) such a Stackelberg payoff is arbitrarily high. In this way, Proposition 3 guarantees that in sequential equilibrium, even in the prescence of reputational incentives, the payoff set can become large absent strong assumptions (stronger than $c^{*}$ incentive compatibility) on the exit function. One consequence of Proposition 3 is that as $\mu_{0}(\omega_{R})\to 1$, the set of sequential equilibria payoffs with reputation converges to the set of sequential equilibria payoffs without, in stark contrast to the baseline payoff selection results of Fudenberg and Levine (1989) without exit.
Proposition 3 highlights the two important roles of Markov equilibrium in establishing the positive reputation result of Theorem 1. First, it disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward tomorrow actions which are not the commitment action. Second, it ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period (c.f. the discussion succeeding Theorem 1. Markov beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today. The proof is relegated to the PROOF OF PROPOSITION 3
4 The Chain Store Game
Consider the canonical chain store game, with two variations: first, suppose the game is infinitely repeated, so the monopolist must continually face entrants over and over again, and second, suppose player actions can lead to exit in the game. The first assumptions implies that, without exit, the results of Fudenberg and Levine (1989) apply, and hence the long-run player can secure their Stackelberg payoff in all Nash equilibria as they become extremely patient. The goal of this section is to understand how exit perturbs these results. In incumbent-entrant games, exit may be a particularly appropriate assumption: accommodating an entrant can lead to the incumbent being displaced form their position, while fighting may increase the probability they remain an incumbent even after entry. This can lead to an additional channel for and challenges to reputation formation.
Formally, suppose in each period the long-run player ($I$) chooses to fight or accomodate an entrant, but only after the entrant ($E$) has decided whether or not to enter. Let stage game payoffs be as follows, where the first entry is the long-run player’s payoff. [^12] These specific payoffs for the chain store game are due to Harry Pei’s Lecture Notes. Interpret here “accomodate” as the action 0.
[…]
Figure 1: Payoffs in the Chain Store Game
Moreover, define $f(a,c)$ to take the values in 1, where $\varphi,\varepsilon,\eta$ are strictly positive and $1-\eta\geq\max\{\varphi,\varepsilon\}$, so survival occurs with higher probability if the entrant chooses Out.
Fight Accomodate
In $\varphi$ $\varepsilon$
Out $1-\eta$ $1-\eta$
Table 1: Exit Probabilities for the Chain Store Game
Note that these payoffs imply the long-run player’s commitment action $c^{*}$ is $F$ for all $\delta$, but that accomodate is the unique Nash equilibrium. These observations (in addition to ranking $F\succ A=0$) imply that Assumption 1 is satisfied. The chain store game admits the following result.
@@ 4 The Chain Store Game
Fix $\varphi>0$. There exists $\overline{\varepsilon}>0$ such that for all (and only) $\varepsilon<\overline{\varepsilon}$
$\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|=0.$
Moreover, there exists $e(\cdot)$, vanishing at 0 such that
$\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq e(\varepsilon).$
The formal proof is in the PROOF OF PROPOSITION 4 The intuition behind the first part is to show that as $\varepsilon$ vanishes, payoffs satisfy $c^{*}$-incentive compatibility; since the game is monotone in the order $In\succ Out$, this implies the second statement. The second part follows by taking $\varepsilon$ small and constructing a revelation equilibrium as in the proof of Proposition 3.
Three observations are in order. First is that in Proposition 4, the key parameter was the probability of exit given that the incumbent entered the game. Intuitively, varying $\varepsilon$ varies the effective discount rate but also the incentives to take different actions; lower $\varepsilon$ increase the “cost” of Nash equilibrium and make reputation formation more attractive to the long-run player, especially relative to fighting. Mathematically, this is evidenced through the likelihood ratio $\frac{\varphi}{\varepsilon}$; as $\varepsilon\to 0$, the likelihood ratio of survival after fighting explodes relative to that after accommodating. Second is the contrast of Proposition 4 with known results from reputation a la Fudenberg and Levine (1989). In particular, the second part of the proposition highlights how reputation can fail with exit even with extremely patient players and high likelihood ratios for survival conditional on the commitment action. This shows that the Markovian refinement is critical for any hope of selecting a unique equilibrium, exactly because it rules out nonstationary equilibrium where the exit cost becomes “sunk” to the long-run player, who chooses their stage-game action conditional on the fact they have already survived today. Finally, I am able to give a unique behavioral prediction in Markov strategies in this game in which the incumbent always fights and short-run players never enter (i.e. deterrence is successful), unlike the indeterminancy of behavioral predictions in reputation games in general (c.f. Pei and Li (2021)).
@@ 5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\overline{c}]$ and moreover suppose $\kappa(\overline{c})>1$, so that the government never chooses the maximally feasible value of punishment. [^13] Note $C=[0,\overline{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(0)$, after which it is sufficient to assume $C=\{0,c^{*}(0)\}$ for the remainder of the application. [^14] This is an Inada-type condition that ensures the government’s problem is always well-defined
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(\theta)=\tau^{2}$.  [^15] In practice, I will consider the case where $\tau^{2}$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfect signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
@@
    The conditional distribution $F_{\theta}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
@@
    The conditional distribution $F_{x}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals (i.e. player $i$’s thought about others’ signals after receiving $x$). The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the state is not in a dominance region.
Assumption 2 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable as the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature. [^16] I am not aware of a global games setting studied in the literature that does not satisfy these assumptions; for example it nests normal-normal information structures.
Having seen their private signals, the dissidents independently and simultaneously choose an action $a_{i}\in\{0,1\}$ in each period, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents defect (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt is successful, is given by $f(A,c)=\mathbb{P}_{\theta}(A\geq\theta)$, and is independent of $c$.
Players’ payoffs are given by
@@
Proposition 5 implies not only that $\mathcal{N}(c)$ is single-valued, but that both the set of states which successfully guard against attacks and the cutoff state are themselves monotone at a cutoff, whose evolution is well-behaved. These properties will be useful in the proof of Theorem 2 and is proven in the PROOF OF PROPOSITION 5 Note Assumption 1 is satisfied as well as $1-\kappa(c)<1-\kappa(0)$ for all $c>0$, and the exit probability depends only on $A$.
Finally, the game is clearly monotone when $A$ is ordered in the standard way in $\mathbb{R}$. Hence the globalized game satisfies the assumptions of the general model.